how yo find velociyt of centre of a bullet w.r.t ground after it stirkes a hoop o negligible thickness with initial vel. vo and getss embedded to it?

Grade:11

1 Answers

Vikas TU
7 years ago
let u be the mass density of the hoop.
mass would be   = u*l where l is the length of the hoop.
Thus from momentum conservation,
m*vo = (m+ u*l)v
v = mvo /(m + ul) m/s would be the final velocity.